# Key Words.: cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD1 Introduction

### v1

@@ 3 Galaxy samples
Our goal is to investigate systematic errors in the measurement of stellar kinematics of massive elliptical galaxies, which comprise the vast majority of deflectors in galaxy-scale strong gravitational lenses. To be sensitive to percent level effects in the systematic errors and covariance, we need spectra with sufficiently high SNR of galaxies that are gravitational lenses – or that are in the appropriate redshift and luminosity range to represent strong gravitational lenses.
Our first sample is a set of massive elliptical galaxies observed with the MUSE spectrograph (Bacon et al. 2010) on VLT, described in Section 3.1. Our second sample is a set of gravitational lens galaxies observed with KCWI on the Keck Telescope, described in Section 3.2. We also consider the high-quality spectrum of the lens galaxy RXJ11311231 observed with JWST-NIRSpec (Jakobsen et al. 2022) (Section 3.3). Even though it is only one galaxy, there are many spectral bins within the galaxy, and thus, one can observe the covariance between the spectra. Furthermore, NIRSpec covers the rest frame CaT region, which is difficult to reach from the ground at our redshifts of interest. Thus, NIRSpec allows us to quantify systematics in a different wavelength region from MUSE and KCWI. Finally, we also study a sample of spectra taken by the Sloan collaboration (Section 3.4). Those generally have a much lower SNR than the other spectra considered here. However, they have been broadly used in the literature. Therefore, we think it might be of interest to the reader to see how our considerations apply to the SDSS dataset.
Throughout this section, we use a SNR metric to compare the quality of data across different instruments. Our chosen metric is the average SNR per Å, measured in the rest frame wavelength interval of 4000–4500 Å (chosen to be redward of the 4000 Å break, where the SNR per pixel is relatively uniform). This is obtained by averaging the signal-to-noise per pixel, divided by the square root of the number of Å per pixel. This metric is not applicable to our JWST data, owing to the spectral range covered. For that dataset, we modify the wavelength range to 8303-8838 Å. For comparison across different wavelength ranges and surveys, it is also interesting to consider the SNR per 70 km s-1 interval (the SDSS pixel resolution). At our fiducial pivot point 4250 Å, 70 km s-1 is almost exactly 1 Å.
@@ 3.1 VLT-MUSE
The Multi-Unit Spectroscopic Explorer (MUSE) is an integral-field spectrograph situated at the Very Large Telescope (VLT) of the European Southern Observatory (ESO). It provides a spatial sampling of 0.2′′across a 1′2 field-of-view (FOV) in its wide-field mode (Bacon et al. 2010). It operates within a wavelength range of 4650-9300 Å with a reciprocal dispersion of 1.25 Å per pixel. The FWHM spectral resolution is approximately 2.5 Å around 5000 Å which corresponds to an instrumental dispersion of 65 km . The galaxies in the original MUSE sample are selected using the integral field unit (IFU) datacubes from the ESO Science Archive under the MUSE-DEEP program 555https://doi.org/10.18727/archive/42 which consolidates observations from single or multiple MUSE programs at specific sky locations into a single IFU datacube, achieving very high effective exposure times. The selection criteria, data extraction process, and other relevant information about the sample, such as galaxy ID and the galaxy cluster or field to which it belongs, ESO archive ID of the datacube, program IDs of the observations used to create the datacube, effective exposure, etc., are described in Mozumdar et al. (2025, in prep; hereafter M25). From the parent sample of M25, we first select a subsample of 140 galaxies that contain data within our adopted restframe wavelength range of 3600-4500 Å with SNR30 Å-1. The central spectra of these galaxies were visually inspected to ensure that no reduction artifacts were present. A handful of spectra were removed due to reduction artifacts, poor sky subtraction, and the presence of strong emission lines. Finally, a velocity dispersion cut of 140 km s-1 was applied to the sample. The reason for this cut is described in Section 4.1. In total, the final sample consists of 85 galaxies with an average integrated SNR of 60 Å-1. The redshift range of this sample is 0.29-0.67. Some examples of the central spectra are shown in Figure 4.
@@ 3.2 Keck-KCWI
Fourteen of the Sloan Lens ACS Survey (SLACS) gravitational lenses (Bolton et al. 2006, 2008) were observed with KCWI on Keck, with the goal of measuring spatially resolved stellar kinematics. The data are fully described by Knabel et al. (2024). The data are collected in the wavelength range 3500-5600 Å in a field of view, and the spectral resolution of R 3600 corresponds to an instrumental dispersion km . The reciprocal dispersion is 0.5 Å per pixel. The redshift range of the sample is 0.15-0.35. In order to attain a high SNR and avoid contamination from the lensed background source, we integrate the datacube within a circular radius of centered on the center of the foreground deflector galaxy. For most of our studies and unless otherwise specified, we use these -integrated spectra. The average SNR of the sample is 160 Å-1 within this radius. For comparison with SDSS spectra of these objects, we also extract spectra from the KCWI datacube integrated within the SDSS aperture-radius of . These spectra are shown in Figure 5 together with the corresponding SDSS spectra for comparison. We use these spectra exclusively for comparison with SDSS. We note that the SDSS spectra for this subsample of objects have insufficient SNR for our purposes and, therefore, will not be used in the overall analysis. We will, however, use a set of higher SNR spectra from SDSS, as described below.
@@ 3.3 JWST/NIRSpec
The gravitational lens RXJ11311231 was observed by JWST/NIRSpec in IFS mode (Böker et al. 2022) during Cycle 1 (Program 1794; PI Suyu, Co-PI Treu). The data analysis and processing are fully described by Shajib et al. (2025; in prep). We constructed the reduced datacube with a spaxel size of 01. The redshift of the lens is . The white-light image of the NIRSpec datacube and one example spectrum is shown in Figure 6.
@@
We take 21 single spaxels within the central region, excluding the four at the corners, of the lens galaxy (illustrated in Figure 6), over which we average our test results. We restrict the bins within the central region of the lens galaxy to minimize contamination from the background quasar and its host galaxy in the spectra adopted for the test. Additionally, we select a wavelength range (8303–8838 Å) sufficiently encompassing the Ca ii triplet lines. This region also includes the H- broad emission + [N ii] lines from the quasar plus host galaxy appearing slightly blueward of the Ca ii triplet and the [S ii] lines from the host in between the two prominent Ca ii lines at 8600 Å. We model all these lines from the quasar and its host with additional kinematic components in our fit (for details, see Shajib et al., 2025, in preparation). However, for the tests performed in this paper, we only focus on the extracted kinematics from the lens galaxy’s Ca ii triplet features.
@@
We remove some spectra from the Indo-US and XSL stellar template libraries that have NaN values or coverage gaps within the fitted wavelength range, leaving us with 609 and 489 templates in them, respectively. We do not include the MILES stellar template library for the tests performed on the NIRSpec data as it does not cover the Ca ii triplet’s wavelength region.
@@
When extracting kinematics from the IFU data, it is standard practice in the literature (e.g., Westfall et al. 2019; Shajib et al. 2023) to first synthesize a single global template666If the spatial extent is sufficient to span population gradients – a set of templates in radial annuli (Knabel et al. 2024). This is not necessary in this case, given that aperture corresponds to just 1 kpc.. This single global template is obtained from the best-fit weighted combination of the whole template library when fitting to the spectra summed from multiple central spaxels or bins, which has a considerably larger SNR. After that, all the individual bins are fit using this single global template. This approach reduces bin-to-bin scatter in the extracted kinematic map for a single galaxy. The creation of a global template from a central region of the galaxy is included as an additional step to the methodology described in the previous section. We find that this additional step indeed reduces bin-to-bin scatter and, thus, in turn, the systematic scatter between the template libraries. Although this is helpful in reducing systematic scatter for the adopted case of NIRSpec spectra, this step may not be necessary for cases with a much higher SNR per spatial bin, where contamination from quasar or host galaxy lines can be robustly differentiated, unlike the specific case presented in this paper.
@@ 3.4 SDSS
For reference and comparison, we also consider SDSS spectra of a sample of galaxies that were identified by Bolton et al. (2008); Shu et al. (2015) as potential gravitational lenses. Targets were then selected from this list for follow-up with the Hubble Space Telescope, and the confirmed lenses represent the SLACS (Bolton et al. 2008) and “SLACS for the masses” (Shu et al. 2015) lenses. Multiple velocity dispersion measurements are available for this sample from the early SDSS releases, the SLACS papers themselves, and the most recent data release, DR17 (Abdurro’uf et al. 2022). We refer to Knabel et al. (2024) for a detailed comparison between the KCWI and SDSS spectra. In general, the SDSS spectra have much lower SNR than required for the present analysis, as shown in Figure 5. However, 50 of them have SNR Å-1. We present results for those here. An example of this set of relatively high SNR SDSS spectra is shown in Figure 7. This subsample of 50 SLACS lens candidates spans a lower redshift range than the KCWI sample () with an average SNR Å-1.
@@ 4 Testing for systematics
In this section, we describe a number of systematic tests conducted on our galaxy samples. As we will see, the dominant source of systematic errors turns out to be the choice of stellar templates, provided that reasonable choices are made for wavelength range and continuum polynomials. We first discuss systematic errors associated with the spectral resolution of the templates in Section 4.1. Then we discuss continuum polynomials (Section 4.2), cleaning templates (Section 4.3), choice of templates (Section 4.4), and wavelength range (Section 4.5).
@@ 4.1 Spectral resolution of the templates
In practice, the stellar velocity dispersion is inferred by comparing the data to a template, initially broadened to match the spectral resolution of the instrumental setup, to find the additional broadening imposed by the stellar kinematics.
Under a Gaussian approximation, the instrumental resolution of the spectrograph used to observe the template , and that used to observe the galaxy spectrum , are related to the total smoothing of the observed spectrum and the astrophysical stellar velocity dispersion by the equation:
For small errors in the characterization of and , it is easy to show – by expansion to leading order – that the relative variations of the astrophysical velocity dispersion and the instrumental and template resolution measurements are related by the following equations. The relative variation in the astrophysical dispersion, at fixed and is:
While the relative variation at fixed and is:
These variations are generally sub-percent. For example, even for a 10% error in the determination of an instrumental resolution of 50 km s-1, the corresponding error on a velocity dispersion of 250 km s-1 is 0.4%. If they are random, e.g., stemming from instabilities of the spectral configuration of slit-filling effects in variable seeing, they will average out for a sample. If they are not random, one should ensure that they are below the noise floor required by the measurement.
For example, the MILES library has a resolution of approximately km s-1at 4000 Å, so if the goal is 1% errors on – assuming that the error on is 3% (Falcón-Barroso et al. 2011) – according to Eq 2 it should not be used for stellar velocity dispersions below 140 km s-1. In contrast, if the goal is 5%, MILES can be used down to 60 km s-1. Since the goal of this paper is 1% we discard galaxies for which the higher resolution libraries (e.g. XSL) give velocity dispersions below 140 km s-1, because MILES would not meet our accuracy standards for this paper and therefore we could not carry out the comparison between libraries.
4.2 Continuum/flux calibration Polynomials
@@
Additive polynomials are less computationally expensive than multiplicative ones and are often preferred for large datasets. However, the computational cost of multiplicative polynomials is still manageable for the sample described here, taking just a few minutes per galaxy. Our extensive tests show that there is usually a “sweet spot” in polynomial orders, where the results from the fit are very stable with respect to specific choices of templates or wavelength ranges. Choosing significantly lower or higher orders can result in significant discrepancies between the results. Therefore, we recommend running a grid of polynomial orders to identify the most appropriate ones for a specific dataset, depending on the wavelength range and accuracy of the spectrophotometric calibration (see, e.g., D’Ago et al. 2023, for a similar strategy).
In practice, we find that additive polynomials of order 6 (but also 5 or 7) and multiplicative polynomials of order 2 (but also 1 or 3) are appropriate for the adopted wavelength range (described in Section 4.5) for our MUSE/KCWI/SDSS datasets. For JWST-NIRSpec, we adopt an additive polynomial order of and a multiplicative polynomial order of , owing to the shorter wavelength range and flatter overall continuum. As we will show in the next section, the residual systematic uncertainty associated with polynomial order choice is at the level of 0.2-0.5% for the datasets considered here.
@@ 4.3 Cleaning the libraries
The effect of cleaning the libraries is significant. Flawed template spectra can cause a variety of issues with the resulting model fit. For example, in the stellar spectrum in the lowest panel of Figure 3, the emission lines in the centers of the CaHK absorption features could artificially broaden and reshape the line profiles in combination with other stellar templates used in the fit. In Figure 8, we show the ratio of velocity dispersion measurements of the MUSE sample using ‘full’777We note that there are some spectra in the full libraries which contain only 0 and NaN. Those need to be excluded to prevent pPXF from crashing. and ‘clean’ XSL, MILES, and Indo-US libraries (the results for the other samples are similar). We see that the effect is of order percent on average for MILES and Indo-US, but it is significantly larger for XSL. In some cases, the differences can be several percent, up to 5% or more. It is thus crucial to use the cleaned libraries for any high-precision measurement.

### v2

@@ 3 Galaxy samples
Our goal is to investigate systematic errors in the measurement of stellar kinematics of massive elliptical galaxies, which comprise the vast majority of deflectors in galaxy-scale strong gravitational lenses. To be sensitive to percent level effects in the systematic errors and covariance, we need spectra with sufficiently high S/N from galaxies that act as gravitational lenses – or that reside within the appropriate redshift and luminosity ranges that typically characterize strong gravitational lenses.
Our first sample is a set of massive elliptical galaxies observed with the MUSE spectrograph (Bacon2010) on VLT, described in Section 3.1. Our second sample is a set of gravitational lens galaxies observed with KCWI on the Keck Telescope, described in Section 3.2. We also consider the high-quality spectrum of the lens galaxy RXJ11311231 observed with JWST-NIRSpec (NIRSpec), as described in Section 3.3. Even though it is only one galaxy, there are many spectral bins within the galaxy, and thus, we can observe the covariance between the spectra. Furthermore, NIRSpec covers the rest frame CaT region, which is difficult to reach from the ground at our redshifts of interest. Thus, NIRSpec allows us to quantify systematics in a different wavelength region than that of MUSE and KCWI. We also included a sample of spectra taken by the Sloan collaboration (Section 3.4). Those generally have a much lower S/N than the other spectra considered here. However, they have been broadly used in the literature. Therefore, we believe it is valuable to demonstrate how our considerations might apply to the SDSS dataset.
Throughout this section, we used a S/N metric to compare the quality of data across different instruments. Our chosen metric is the average S/N per Å, measured in the rest frame wavelength interval of 4000–4500 Å (chosen to be redward of the 4000 Å break, where the S/N per pixel is relatively uniform). This is obtained by averaging the S/N per pixel, divided by the square root of the number of Å per pixel. This metric is not applicable to our JWST data, owing to the spectral range covered. For that dataset, we modify the wavelength range to 8303-8838 Å. When making comparisons across different wavelength ranges and surveys, it is also interesting to consider the S/N per 70 km s-1 interval (the SDSS pixel resolution). At our fiducial pivot point 4250 Å, 70 km s-1 is almost exactly 1 Å.
@@ 3.1 VLT-MUSE
The Multi-Unit Spectroscopic Explorer (MUSE) is an integral-field spectrograph situated at the Very Large Telescope (VLT) of the European Southern Observatory (ESO). It provides a spatial sampling of 0.2′′across a 1′2 field of view (FoV) in its wide-field mode (Bacon2010). It operates within a wavelength range of 4650-9300 Å with a reciprocal dispersion of 1.25 Å per pixel. The FWHM spectral resolution is approximately 2.5 Å around 5000 Å which corresponds to an instrumental dispersion of 65 km . The galaxies in the original MUSE sample are selected using the integral field unit (IFU) datacubes from the ESO Science Archive under the MUSE-DEEP program 777https://doi.org/10.18727/archive/42, which consolidates observations from single or multiple MUSE programs at specific sky locations into a single IFU datacube, achieving very high effective exposure times. The selection criteria, data extraction process, and other relevant information about the sample, such as the galaxy ID and the galaxy cluster or field to which it belongs, ESO archive ID of the datacube, program IDs of the observations used to create the datacube, effective exposure, and other details are described in MAGNUS_I. From the parent sample of MAGNUS_I, we first selected a subsample of 140 galaxies that contain data within our adopted restframe wavelength range of 3600-4500 Å with S/N30 Å-1. The central spectra of these galaxies were visually inspected to ensure that no reduction artifacts were present. A handful of spectra were removed due to reduction artifacts, poor sky subtraction, and the presence of strong emission lines. Finally, a velocity dispersion cut of 140 km s-1 was applied to the sample. The reason for this cut is described in Section 4.1. In total, the final sample consists of 85 galaxies with an average integrated S/N of 60 Å-1. The redshift range of this sample is 0.29-0.67. Some examples of the central spectra are shown in Figure 4.
@@ 3.2 Keck-KCWI
Fourteen of the Sloan Lens ACS Survey (SLACS) gravitational lenses (Bolton06; Bolton08) were observed with KCWI on Keck, with the goal of measuring spatially resolved stellar kinematics. The data are fully described by Knabel24. The data are collected in the wavelength range 3500-5600 Å in a FoV, and the spectral resolution of R 3600 corresponds to an instrumental dispersion km . The reciprocal dispersion is 0.5 Å per pixel. The redshift range of the sample is 0.15-0.35. In order to attain a high S/N and avoid contamination from the lensed background source, we integrated the datacube within a circular radius of centered on the center of the foreground deflector galaxy. For most of our studies (and unless otherwise specified), we used these -integrated spectra. The average S/N of the sample is 160 Å-1 within this radius. For comparison with SDSS spectra of these objects, we also extracted spectra from the KCWI datacube integrated within the SDSS aperture-radius of . These spectra are shown in Figure 5 together with the corresponding SDSS spectra for comparison. We use these spectra exclusively for comparison with SDSS. We note that the SDSS spectra for this subsample of objects have insufficient S/Ns for our purposes and were therefore not used in the overall analysis. We did, however, use a set of higher S/N spectra from SDSS, as described below.
@@ 3.3 JWST/NIRSpec
The gravitational lens RXJ11311231 was observed by JWST/NIRSpec in IFS mode (NIRSpec22) during Cycle 1 (Program 1794; PI Suyu, Co-PI Treu). The data analysis and processing are fully described by Shajib25. We constructed the reduced datacube with a spaxel size of 01. The redshift of the lens is . The white-light image of the NIRSpec datacube and one example spectrum is shown in Figure 6.
@@
We took 21 single spaxels within the central region, excluding the four at the corners, of the lens galaxy (illustrated in Figure 6), over which we average our test results. We restricted the bins within the central region of the lens galaxy to minimize contamination from the background quasar and its host galaxy in the spectra adopted for the test. Additionally, we selected a wavelength range (8303–8838 Å) sufficiently encompassing the Ca ii triplet lines. This region also includes the H broad emission + [N ii] lines from the quasar plus host galaxy appearing slightly blueward of the Ca ii triplet and the [S ii] lines from the host in between the two prominent Ca ii lines at 8600 Å. We modeled all these lines from the quasar and its host with additional kinematic components in our fit (for details, see Shajib25). However, for the tests performed in this paper, we only focused on the extracted kinematics from the lens galaxy’s Ca ii triplet features.
@@
We removed some spectra from the Indo-US and XSL stellar template libraries that have NaN values or coverage gaps within the fitted wavelength range, leaving us with 609 and 489 templates in them, respectively. We did not include the MILES stellar template library for the tests performed on the NIRSpec data as it does not cover the Ca ii triplet’s wavelength region.
@@
When extracting kinematics from the IFU data, it is standard practice in the literature (e.g., Manga19; Shajib23) to first synthesize a single global template. This single global template is obtained from the best-fit weighted combination of the whole template library when fitting to the spectra summed from multiple central spaxels or bins, which has a considerably larger S/N. Afterward, all the individual bins are fit using this single global template. This approach reduces bin-to-bin scatter in the extracted kinematic map for a single galaxy. The creation of a global template from a central region of the galaxy is included as an additional step to the methodology described in the previous section. We find that this additional step indeed reduces bin-to-bin scatter and, thus, in turn, the systematic scatter between the template libraries. Although this is helpful in reducing systematic scatter for the adopted case of NIRSpec spectra, this step might not be necessary for cases with a a much higher S/N per spatial bin, where contamination from quasar or host galaxy lines can be robustly differentiated, unlike the specific case presented in this paper.
@@ 3.4 SDSS
For reference and comparison, we also consider SDSS spectra of a sample of galaxies that were identified by Bolton08; Shu15 as potential gravitational lenses. Targets were then selected from this list for follow-up with the Hubble Space Telescope, and the confirmed lenses represent the SLACS (Bolton08) and “SLACS for the masses” (Shu15) lenses. Multiple velocity dispersion measurements are available for this sample from the early SDSS releases, the SLACS papers themselves, and the most recent data release, DR17 (DR17). We refer to Knabel24 for a detailed comparison between the KCWI and SDSS spectra. In general, the SDSS spectra have much lower S/N than required for the present analysis, as shown in Figure 5. However, 50 of them have S/N Å-1. We present results for those here. An example of this set of relatively high-S/N SDSS spectra is shown in Figure 7. This subsample of 50 SLACS lens candidates spans a lower redshift range than the KCWI sample () with an average S/N Å-1.
@@ 4 Testing for systematics
In this section, we describe a number of systematic tests conducted on our galaxy samples. As we explain here, the dominant source of systematic errors turns out to be the choice of stellar templates, provided that reasonable choices are made for wavelength range and continuum polynomials. We first discuss systematic errors associated with the spectral resolution of the templates in Section 4.1. Then we discuss continuum polynomials (Section 4.2), cleaning templates (Section 4.3), choice of templates (Section 4.4), and wavelength range (Section 4.5).
@@ 4.1 Spectral resolution of the templates
In practice, the stellar velocity dispersion is inferred by comparing the data to a template. If the template resolution is matched to the data, then the astrophysical is the one returned by ppxf without the need for changes. Templates of higher resolution than the data are initially broadened to match the spectral resolution of the instrumental setup, to find the additional broadening imposed by the stellar kinematics. For templates with lower resolution than the observed data, the effect of this broadening must be applied as a correction after measuring the dispersion from the data.
Under a Gaussian approximation, the instrumental resolution of the spectrograph used to observe the template , and that used to observe the galaxy spectrum , are related to the total smoothing of the observed spectrum and the astrophysical stellar velocity dispersion via the equation
For small errors in the characterization of and , it is easy to show – by expansion to leading order – that the relative variations of the astrophysical velocity dispersion and the instrumental and template resolution measurements are related by the following equations. The relative variation in the astrophysical dispersion, at fixed and is,
While the relative variation at fixed and is
These variations are generally sub-percent. For example, even for a 10% error in the determination of an instrumental resolution of 50 km s-1, the corresponding error on a velocity dispersion of 250 km s-1 is 0.4%. If they are random (e.g., stemming from instabilities of the spectral configuration of slit-filling effects in variable seeing), they will average out for a sample. If they are not random, one should ensure that they are below the noise floor required by the measurement.
For example, the MILES library has a resolution of approximately km s-1at 4000 Å; thus, if the goal is 1% errors on , and assuming that the error on is 3% (Falcon-Barroso11), and according to Eq, 2, it should not be used for stellar velocity dispersions below 140 km s-1. In contrast, if the goal is 5%, MILES can be used down to 60 km s-1. Since the goal of this paper is 1% we discarded galaxies for which the higher resolution libraries (e.g. XSL) give velocity dispersions below 140 km s-1, because MILES would not meet our standards of precision for this paper and therefore we would not be able to carry out the comparison between libraries.
4.2 Continuum and flux calibration polynomials
@@
Additive polynomials are less computationally expensive than multiplicative ones and are often preferred for large datasets. However, the computational cost of multiplicative polynomials is still manageable for the sample described here, taking just a few minutes per galaxy. Our extensive tests show that there is usually a “sweet spot” in polynomial orders, where the results from the fit are very stable with respect to specific choices of templates or wavelength ranges. Choosing significantly lower or higher orders can result in significant discrepancies between the results. Therefore, we recommend running a grid of polynomial orders to identify the most appropriate ones for a specific dataset, depending on the wavelength range and accuracy of the spectrophotometric calibration (see, e.g., DAGO23, for a similar strategy).
In practice, we find that additive polynomials of order 6 (as well as 5 or 7) and multiplicative polynomials of order of 2 (as well as 1 or 3) are appropriate for the adopted wavelength range (described in Section 4.5) for our MUSE/KCWI/SDSS datasets. For JWST-NIRSpec, we adopted an additive polynomial order of and a multiplicative polynomial order of , owing to the shorter wavelength range and flatter overall continuum. As we show in the next section, the residual systematic uncertainty associated with polynomial order choice is at the level of 0.2-0.5% for the datasets considered here.
@@ 4.3 Cleaning the libraries
The effect of cleaning the libraries is significant. Flawed template spectra can cause a variety of issues with the resulting model fit. For example, in the stellar spectrum in the lowest panel of Figure 2, the emission lines in the centers of the CaHK absorption features could artificially broaden and reshape the line profiles in combination with other stellar templates used in the fit. In Figure 8, we show the ratio of velocity dispersion measurements of the MUSE sample using ‘full’888We note that there are some spectra in the full libraries which contain only 0 and NaN. Those need to be excluded to prevent pPXF from crashing. and ‘clean’ XSL, MILES, and Indo-US libraries (the results for the other samples are similar). We see that the effect is of order percent on average for MILES and Indo-US, but it is significantly larger for XSL. In some cases, the differences can be several percent, up to 5% or more. It is thus crucial to use the cleaned libraries for any high-precision measurement.